solve each system using the substitution method.
s+t=-5
s-t=3

Answer by london maths tutor(243)

s+t=-5 ... equation 1
s-t=3 ... equation 2
Rearrange equation 2:
s-t = 3
s = t + 3
Put s = t+3 into equation 1
(t+3)+t = -5
3 + 2t = -5
2t = -8
t = -4
Put t = -4 into equation 2
s -(-4) = 3
s + 4 = 3
s = -1
Answer: s = -1 and t = -4